# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ III-A1 General Purpose
A causal decoder trained on the same dataset as the Gopher [111] but with a little different data sampling distribution (sampled from MassiveText). The model architecture is similar to the one used for Gopher, with the exception of AdamW optimizer instead of Adam. Chinchilla identifies the relationship that model size should be doubled for every doubling of training tokens. Over 400 language models ranging from 70 million to over 16 billion parameters on 5 to 500 billion tokens are trained to get the estimates for compute-optimal training under a given budget. The authors train a 70B model with the same compute budget as Gopher (280B) but with 4 times more data. It outperforms Gopher [111], GPT-3 [6], and others on various downstream tasks, after fine-tuning.
AlexaTM [120]
@@
An encoder-decoder model, where encoder weights and decoder embeddings are initialized with a pre-trained encoder to speedup training. The encoder stays frozen for initial 100k steps and later unfreezed for end-to-end training. The model is trained on a combination of denoising and causal language modeling (CLM) objectives, concatenating $[CLM]$ token at the beginning for mode switiching. During training, the CLM task is applied for 20% of the time, which improves the in-context learning performance.
PaLM [15]
@@
A causal decoder with parallel attention and feed-forward layers similar to Eq. 4, speeding up training 15 times faster. Additional changes to the conventional transformer model include SwiGLU activation, RoPE embeddings, multi-query attention that saves computation cost during decoding, and shared input-output embeddings. During training, loss spiking was observed, and to fix it, model training was restarted from a 100 steps earlier checkpoint by skipping 200-500 batches around the spike. Moreover, the model was found to memorize around 2.4% of the training data at the 540B model scale, whereas this number was lower for smaller models.
@@
  PaLM-2 [121]: A smaller multi-lingual variant of PaLM, trained for larger iterations on a better quality dataset. The PaLM-2 shows significant improvements over PaLM, while reducing training and inference costs due to its smaller size. To lessen toxicity and memorization, it appends special tokens with a fraction of pre-training data, which shows reduction in generating harmful responses.
U-PaLM [122]
@@
This method trains PaLM for 0.1% additional compute with UL2 (also named as UL2Restore) objective [123] using the same dataset and outperforms baseline significantly on various NLP tasks, including zero-shot, few-shot, commonsense reasoning, CoT, etc. Training with UL2R involves converting a causal decoder PaLM to a non-causal decoder PaLM and employing 50% sequential denoising, 25% regular denoising, and 25% extreme denoising loss functions.
UL2 [123]
@@
An encoder-decoder architecture trained using a mixture of denoisers (MoD) objectives. Denoisers include 1) R-Denoiser: a regular span masking, 2) S-Denoiser: which corrupts consecutive tokens of a large sequence and 3) X-Denoiser: which corrupts a large number of tokens randomly. During pre-training, UL2 includes a denoiser token from ${R,S,X}$ to represent a denoising setup. It helps improve fine-tuning performance for downstream tasks that bind the task to one of the upstream training modes. This MoD style of training outperforms the T5 model on many benchmarks.
GLM-130B [33]
@@
GLM-130B is a bilingual (English and Chinese) model trained using an auto-regressive mask infilling pre-training objective similar to the GLM [124]. This training style makes the model bidirectional as compared to GPT-3, which is unidirectional. Opposite to the GLM, the training of GLM-130B includes a small amount of multi-task instruction pre-training data (5% of the total data) along with the self-supervised mask infilling. To stabilize the training, it applies embedding layer gradient shrink.
LLaMA [125, 21]
@@
A set of decoder-only language models varying from 7B to 70B parameters. LLaMA models series is the most famous among the community for parameter-efficient and instruction tuning.
@@
  LLaMA-1 [125]: Implements efficient causal attention [126] by not storing and computing masked attention weights and key/query scores. Another optimization is reducing number of activations recomputed in backward pass, as in  [127].
@@
  LLaMA-2 [21]: This work is more focused towards fine-tuning a safer and better LLaMA-2-Chat model for dialogue generation. The pre-trained model has 40% more training data with a larger context length and grouped-query attention.
PanGu-$\Sigma$ [128]
@@
An autoregressive model with parameters copied from PanGu-$\alpha$ and extended to a trillion scale with Random Routed Experts (RRE), the architectural diagram is shown in Figure 10. RRE is similar to the MoE architecture, with distinctions at the second level, where tokens are randomly routed to experts in a domain instead of using a learnable gating method. The model has bottom layers densely activated and shared across all domains, whereas top layers are sparsely activated according to the domain. This training style allows extracting task-specific models and reduces catastrophic forgetting effects in case of continual learning.
III-A2 Coding
CodeGen [129]
@@ III-A2 Coding
CodeGen has similar architecture to the PaLM [15], i.e., parallel attention, MLP layers, and RoPE embeddings. The model is trained on both natural language and programming language data sequentially (trained on the first dataset, then the second and so on) on the following datasets 1) PILE, 2) BIGQUERY and 3) BIGPYTHON. CodeGen proposed a multi-step approach to synthesizing code. The purpose is to simplify the generation of long sequences where the previous prompt and generated code are given as input with the next prompt to generate the next code sequence. CodeGen opensource a Multi-Turn Programming Benchmark (MTPB) to evaluate multi-step program synthesis.
Codex [130]
@@
This LLM is trained on a subset of public Python Github repositories to generate code from docstrings. Computer programming is an iterative process where the programs are often debugged and updated before fulfilling the requirements. Similarly to this, Codex generates 100 versions of a program by repetitive sampling for a given description, which produces a working solution for 77.5% of the problems passing unit tests. Its powerful version powers Github Copilot [^2] https://github.com/features/copilot.
AlphaCode [131]
@@
A set of large language models, ranging from 300M to 41B parameters, designed for competition-level code generation tasks. It uses the multi-query attention [132] to reduce memory and cache costs. Since competitive programming problems highly require deep reasoning and an understanding of complex natural language algorithms, the AlphaCode models are pre-trained on filtered GitHub code in popular languages and then fine-tuned on a new competitive programming dataset named CodeContests. The CodeContests dataset mainly contains problems, solutions, and test cases collected from the Codeforces platform [^3] https://codeforces.com/. The pre-training employs standard language modeling objectives, while GOLD [133] with tempering [134] serves as the training objective for the fine-tuning on CodeContests data. To evaluate the performance of AlphaCode, simulated programming competitions are hosted on the Codeforces platform: overall, AlphaCode ranks at the top 54.3% among over 5000 competitors, where its Codeforces rating is within the top 28% of recently participated users.
CodeT5+ [34]
@@
CodeT5+ is based on CodeT5 [135], with shallow encoder and deep decoder, trained in multiple stages initially unimodal data (code) and later bimodal data (text-code pairs). Each training stage has different training objectives and activates different model blocks encoder, decoder, or both according to the task. The unimodal pre-training includes span denoising and CLM objectives, whereas bimodal pre-training objectives contain contrastive learning, matching, and CLM for text-code pairs. CodeT5+ adds special tokens with the text to enable task modes, for example, $[CLS]$ for contrastive loss, $[Match]$ for text-code matching, etc.
StarCoder [136]
@@
A decoder-only model with SantaCoder architecture, employing Flash attention to scale up the context length to 8k. The StarCoder trains an encoder to filter names, emails, and other personal data from the training data. Its fine-tuned variant outperforms PaLM, LLaMA, and LAMDA on HumanEval and MBPP benchmarks.
III-A3 Scientific Knowledge
Galactica [137]
@@ III-A3 Scientific Knowledge
A large curated corpus of human scientific knowledge with 48 million papers, textbooks, lecture notes, millions of compounds and proteins, scientific websites, encyclopedias, and more are trained using metaseq library3, which is built on PyTorch and fairscale [138]. The model wraps reasoning datasets with $<work>$ token to provide step-by-step reasoning context to the model, which has been shown to improve the performance on reasoning tasks.
III-A4 Dialog
LaMDA [139]
@@ III-A4 Dialog
A decoder-only model pre-trained on public dialog data, public dialog utterances, and public web documents, where more than 90% of the pre-training data is in English. LaMDA is trained with the objective of producing responses that exhibit high levels of quality, safety, and groundedness. To achieve this, discriminative and generative fine-tuning techniques are incorporated to enhance the model’s safety and quality aspects. As a result, the LaMDA models can be utilized as a general language model performing various tasks.
III-A5 Finance
BloombergGPT [140]
@@ III-A5 Finance
A non-causal decoder model trained using both financial ("FINPILE" from the Bloomberg archive) and general-purpose datasets. The model’s architecture is similar to the BLOOM [13] and OPT [14]. It allocates 50B parameters to different blocks of the model using the approach [108]. For effective training, BloombergGPT packs documents together with $<|endoftext|>$ to use maximum sequence length, use warmup batch size starting from 1024 to 2048, and manually reduces the learning rate multiple times during the training.
Xuan Yuan 2.0 [141]
@@
A Chinese financial chat model with BLOOM’s [13] architecture trained on a combination of general purpose, financial, general purpose instructions, and financial institutions datasets. Xuan Yuan 2.0 combined the pre-training and fine-tuning stages to avoid catastrophic forgetting.
@@
TABLE I: Noteworthy findings and insights from pre-trained Large Language Model.
Models Findings & Insights
@@
Encoder and decoder with shared parameters perform equivalently when parameters are not shared Fine-tuning model layers (adapter layers) work better than the conventional way of training on only classification layers
@@
Few-shot performance of LLMs is better than the zero-shot, suggesting that LLMs are meta-learners
@@
Large multi-lingual models perform equivalently to single language models on downstream tasks. However, smaller multi-lingual models perform worse
@@
LLMs are good at a few shot capabilities
@@
Prompt fine-tuning requires updating very few parameters while achieving performance comparable to full model fine-tuning Prompt fine-tuning takes more time to converge as compared to full model fine-tuning Inserting prompt tokens in-between sentences can allow the model to understand relations between sentences and long sequences In an analysis, CPM-2 finds that prompts work as a provider (additional context) and aggregator (aggregate information with the input text) for the model
Codex
This LLM focuses on code evaluations and introduces a novel way of selecting the best code samples. The results indicate it is possible to accurately select code samples using heuristic ranking in lieu of a detailed evaluation of each sample, which may not be feasible or feasible in some situations.
@@
ERNIE 3.0 shows that a modular LLM architecture with a universal representation module and task-specific representation module helps in finetuning phase. Optimizing the parameters of a task-specific representation network during the fine-tuning phase is an efficient way to take advantage of the powerful pretrained model.
@@
The performance of an LLM is highly related to the network size. To improve runtime performance, more operations can be performed in parallel (width) rather than sequentially (depth). To efficiently represent and fit more text in the same context length, the model uses a larger vocabulary to train a SentencePiece tokenizer without restricting it to word boundaries. This tokenizer improvement can further benefit few-shot learning tasks.
@@
By employing prompt-based tuning, the performances of models can be improved, often surpassing those of state-of-the-art models when the backward gradients of inputs are accessible.
@@
The model architecture that excels in pre-training and fine-tuning cases may exhibit contrasting behavior in zero-shot and few-shot learning.
@@
Relative encodings enable models to be evaluated for longer sequences than those on which it was trained.
@@
This LLM builds on top of ERNIE 3.0 and add a self-supervised adversarial loss to distinguish whether a text is generated or the original one. This distinction ability between real and generate text improves the LLM’s performance as compared to ERNIE 3.0.
@@
Parallel attention + FF layers speed-up training 15% with the same performance as with cascaded layers Initializing feed-forward output layers before residuals with scheme in [142] avoids activations from growing with increasing depth and width Training on Pile outperforms GPT-3 on five-shot
@@
Restart training from an earlier checkpoint with a lower learning rate if loss diverges Model is prone to generate repetitive text and stuck in a loop
BLOOM
None
@@
Galactica’s performance has continued to improve across validation set, in-domain, and out-of-domain benchmarks, even with multiple repetitions of the corpus, which is superior to existing research on LLMs. A working memory token approach can achieve strong performance over existing methods on mathematical MMLU and MATH benchmarks. It sets a new state-of-the-art on several downstream tasks such as PubMedQA (77.6%) and MedMCQA dev (52.9%).
@@
The feed-forward component of each Transformer layer can be replaced with a mixture-of-experts (MoE) module consisting of a set of independent feed-forward networks (i.e., the ‘experts’). By sparsely activating these experts, the model capacity can be maintained while much computation is saved. By leveraging sparsity, we can make significant strides toward developing high-quality NLP models while simultaneously reducing energy consumption. Consequently, MoE emerges as a robust candidate for future scaling endeavors. The model trained on filtered data shows consistently better performances on both NLG and NLU tasks, where the effect of filtering is more significant on the former tasks. Filtered pretraining corpora plays a crucial role in the generation capability of LLMs, especially for the downstream tasks. The scaling of GLaM MoE models can be achieved by increasing the size or number of experts in the MoE layer. Given a fixed budget of computation, more experts contribute to better predictions.
@@
The model can be fine-tuned to learn to call different external information resources and tools.

### 09 Apr 2024 v9

@@ 3.1.1 General Purpose
 Chinchilla [96]: A causal decoder trained on the same dataset as the Gopher [116] but with a little different data sampling distribution (sampled from MassiveText). The model architecture is similar to the one used for Gopher, with the exception of AdamW optimizer instead of Adam. Chinchilla identifies the relationship that model size should be doubled for every doubling of training tokens. Over 400 language models ranging from 70 million to over 16 billion parameters on 5 to 500 billion tokens are trained to get the estimates for compute-optimal training under a given budget. The authors train a 70B model with the same compute budget as Gopher (280B) but with 4 times more data. It outperforms Gopher [116], GPT-3 [6], and others on various downstream tasks, after fine-tuning.
@@
 AlexaTM [122]: An encoder-decoder model, where encoder weights and decoder embeddings are initialized with a pre-trained encoder to speed up training. The encoder stays frozen for the initial 100k steps and is later unfrozen for end-to-end training. The model is trained on a combination of denoising and causal language modeling (CLM) objectives, concatenating a $[CLM]$ token at the beginning for mode switching. During training, the CLM task is applied for 20% of the time, which improves the in-context learning performance.
@@
 PaLM [15]: A causal decoder with parallel attention and feed-forward layers similar to Eq. 4, speeding up training by a factor of 15. Additional changes to the conventional transformer model include SwiGLU activation, RoPE embeddings, multi-query attention that saves computation cost during decoding, and shared input-output embeddings. During training, loss spiking was observed, and to fix it, model training was restarted from a 100-step earlier checkpoint by skipping 200-500 batches around the spike. Moreover, the model was found to memorize around 2.4% of the training data at the 540B model scale, whereas this number was lower for smaller models.
@@
 PaLM-2 [123]: A smaller multi-lingual variant of PaLM, trained for larger iterations on a better quality dataset. PaLM-2 shows significant improvements over PaLM, while reducing training and inference costs due to its smaller size. To lessen toxicity and memorization, it appends special tokens with a fraction of pre-training data, which shows a reduction in generating harmful responses.
@@
 U-PaLM [124]: This method trains PaLM for 0.1% additional compute with the UL2 (also named as UL2Restore) objective [125], using the same dataset it outperforms the baseline significantly on various NLP tasks, including zero-shot, few-shot, commonsense reasoning, CoT, etc. Training with UL2R involves converting a causal decoder PaLM to a non-causal decoder PaLM and employing 50% sequential denoising, 25% regular denoising, and 25% extreme denoising loss functions.
@@
 UL2 [125]: An encoder-decoder architecture trained using a mixture of denoisers (MoD) objective. Denoisers include 1) R-Denoiser: a regular span masking, 2) S-Denoiser: which corrupts consecutive tokens of a large sequence and 3) X-Denoiser: which corrupts a large number of tokens randomly. During pre-training, UL2 includes a denoiser token from R,S,X{R,S,X} to represent a denoising setup. It helps improve fine-tuning performance for downstream tasks that bind the task to one of the upstream training modes. This MoD style of training outperforms the T5 model on many benchmarks.
@@
 GLM-130B [33]: GLM-130B is a bilingual (English and Chinese) model trained using an auto-regressive mask infilling pre-training objective similar to the GLM [126]. This training style makes the model bidirectional as compared to GPT-3, which is unidirectional. As opposed to GLM, the training of GLM-130B includes a small amount of multi-task instruction pre-training data (5% of the total data) along with self-supervised mask infilling. To stabilize the training, it applies embedding layer gradient shrink.
@@
 LLaMA [127, 21]: A set of decoder-only language models varying from 7B to 70B parameters. LLaMA models series is the most famous among the community for parameter efficiency and instruction tuning.
@@
 LLaMA-1 [127]: Implements efficient causal attention [128] by not storing and computing masked attention weights and key/query scores. Another optimization is reducing the number of activations recomputed in the backward pass, as in  [129].
@@
 LLaMA-2 [21]: This work is more focused on fine-tuning a safer and better LLaMA-2-Chat model for dialogue generation. The pre-trained model has 40% more training data with a larger context length and grouped-query attention.
@@
 PanGu-$\Sigma$ [92]: An autoregressive model with parameters copied from PanGu-$\alpha$ and extended to a trillion scale with Random Routed Experts (RRE), the architectural diagram is shown in Figure 10. RRE is similar to the MoE architecture, with distinctions at the second level, where tokens are randomly routed to experts in a domain instead of using a learnable gating method. The model has bottom layers densely activated and shared across all domains, whereas top layers are sparsely activated according to the domain. This training style allows extracting task-specific models and reduces catastrophic forgetting effects in the case of continual learning.
3.1.2 Coding
@@
 CodeGen [130]: CodeGen has similar architecture to PaLM [15], i.e., parallel attention, MLP layers, and RoPE embeddings. The model is trained on both natural language and programming language data sequentially (trained on the first dataset, then the second and so on) on the following datasets 1) PILE, 2) BIGQUERY and 3) BIGPYTHON. CodeGen proposed a multi-step approach to synthesizing code. The purpose is to simplify the generation of long sequences where the previous prompt and generated code are given as input with the next prompt to generate the next code sequence. CodeGen opensource a Multi-Turn Programming Benchmark (MTPB) to evaluate multi-step program synthesis.
@@
 Codex [131]: This LLM is trained on a subset of public Python Github repositories to generate code from docstrings. Computer programming is an iterative process where the programs are often debugged and updated before fulfilling the requirements. Similarly to this, Codex generates 100 versions of a program by repetitive sampling for a given description, which produces a working solution for 77.5% of the problems passing unit tests. Its powerful version powers Github Copilot [^2] https://github.com/features/copilot.
@@
 AlphaCode [132]: A set of large language models, ranging from 300M to 41B parameters, designed for competition-level code generation tasks. It uses the multi-query attention [133] to reduce memory and cache costs. Since competitive programming problems highly require deep reasoning and an understanding of complex natural language algorithms, the AlphaCode models are pre-trained on filtered GitHub code in popular languages and then fine-tuned on a new competitive programming dataset named CodeContests. The CodeContests dataset mainly contains problems, solutions, and test cases collected from the Codeforces platform [^3] https://codeforces.com/. The pre-training employs standard language modeling objectives, while GOLD [134] with tempering [135] serves as the training objective for the fine-tuning on CodeContests data. To evaluate the performance of AlphaCode, simulated programming competitions are hosted on the Codeforces platform: overall, AlphaCode ranks at the top 54.3% among over 5000 competitors, where its Codeforces rating is within the top 28% of recently participated users.
@@
 CodeT5+ [34]: CodeT5+ is based on CodeT5 [136], with shallow encoder and deep decoder, trained in multiple stages initially unimodal data (code) and later bimodal data (text-code pairs). Each training stage has different training objectives and activates different model blocks encoder, decoder, or both according to the task. The unimodal pre-training includes span denoising and CLM objectives, whereas bimodal pre-training objectives contain contrastive learning, matching, and CLM for text-code pairs. CodeT5+ adds special tokens with the text to enable task modes, for example, $[CLS]$ for contrastive loss, $[Match]$ for text-code matching, etc.
@@
 StarCoder [137]: A decoder-only model with the SantaCoder architecture, employing Flash attention to scale up the context length to 8k. The StarCoder trains an encoder to filter names, emails, and other personal data from the training data. Its fine-tuned variant outperforms PaLM, LLaMA, and LAMDA on HumanEval and MBPP benchmarks.
3.1.3 Scientific Knowledge
@@
 Galactica [138]: A large curated corpus of human scientific knowledge with 48 million papers, textbooks, lecture notes, millions of compounds and proteins, scientific websites, encyclopedias, and more are trained using the metaseq library3, which is built on PyTorch and fairscale [139]. The model wraps reasoning datasets with the <work><work> token to provide step-by-step reasoning context to the model, which has been shown to improve the performance on reasoning tasks.
3.1.4 Dialog
@@
 LaMDA [140]: A decoder-only model pre-trained on public dialog data, public dialog utterances, and public web documents, where more than 90% of the pre-training data is in English. LaMDA is trained with the objective of producing responses that exhibit high levels of quality, safety, and groundedness. To achieve this, discriminative and generative fine-tuning techniques are incorporated to enhance the model’s safety and quality aspects. As a result, the LaMDA models can be utilized as a general language model performing various tasks.
3.1.5 Finance
@@
 BloombergGPT [141]: A non-causal decoder model trained using both financial ("FINPILE" from the Bloomberg archive) and general-purpose datasets. The model’s architecture is similar to the BLOOM [13] and OPT [14]. It allocates 50B parameters to different blocks of the model using the approach [113]. For effective training, BloombergGPT packs documents together with <|endoftext|><|endoftext|> to use the maximum sequence length, uses warmup batch size starting from 1024 to 2048, and manually reduces the learning rate multiple times during the training.
@@
 Xuan Yuan 2.0 [142]: A Chinese financial chat model with BLOOM’s [13] architecture trained on a combination of general purpose, financial, general purpose instructions, and financial institutions datasets. Xuan Yuan 2.0 combined the pre-training and fine-tuning stages to avoid catastrophic forgetting.
@@
Table 1: Noteworthy findings and insights of pre-trained Large Language Models.
[…]
@@
1. Encoder and decoder with shared parameters perform equivalently when parameters are not shared 2. Fine-tuning model layers (adapter layers) work better than the conventional way of training on only classification layers
@@
1. Few-shot performance of LLMs is better than the zero-shot, suggesting that LLMs are meta-learners
@@
1. Large multi-lingual models perform equivalently to single language models on downstream tasks. However, smaller multi-lingual models perform worse
@@
1. LLMs have good few shot capabilities
@@
1. Prompt fine-tuning requires updating very few parameters while achieving performance comparable to full model fine-tuning 2. Prompt fine-tuning takes more time to converge as compared to full model fine-tuning 3. Inserting prompt tokens in-between sentences can allow the model to understand relations between sentences and long sequences 4. In an analysis, CPM-2 finds that prompts work as a provider (additional context) and aggregator (aggregate information with the input text) for the model
@@
1. A modular LLM architecture with a universal representation module and task-specific representation module helps in the finetuning phase 2. Optimizing the parameters of a task-specific representation network during the fine-tuning phase is an efficient way to take advantage of the powerful pre-trained model
@@
1. The performance of LLM is highly related to the network size 2. To improve runtime performance, more operations can be performed in parallel (width) rather than sequential (depth) 3. To efficiently represent and fit more text in the same context length, the model uses a larger vocabulary to train a SentencePiece tokenizer without restricting it to word boundaries. This further benefits in few-shot learning tasks
@@
1. By employing prompt-based tuning, the performances of models can be improved, often surpassing those of state-of-the-art models when the backward gradients of inputs are accessible
@@
1. The model architecture that excels in pre-training and fine-tuning cases may exhibit contrasting behavior in zero-shot and few-shot learning
@@
1. Relative encodings enable the model to evaluate for longer sequences than training.
@@
1. Additional self-supervised adversarial loss to distinguish between real and generated text improves the model performance as compared to ERNIE 3.0
@@
1. Parallel attention + $F$ layers speed-up training 15% with the same performance as with cascaded layers 2. Initializing feed-forward output layers before residuals with scheme in [143] avoids activations from growing with increasing depth and width 3. Training on Pile outperforms GPT-3 on five-shot
Table Continued on Next Page
[…]
@@
1. Restart training from an earlier checkpoint with a lower learning rate if loss diverges 2. Model is prone to generate repetitive text and stuck in a loop
@@
1. Galactica’s performance has continued to improve across validation set, in-domain, and out-of-domain benchmarks, even with multiple repetitions of the corpus, which is superior to existing research on LLMs 2. A working memory token approach can achieve strong performance over existing methods on mathematical MMLU and MATH benchmarks. It sets a new state-of-the-art on several downstream tasks such as PubMedQA (77.6%) and MedMCQA dev (52.9%)
@@
1. The model capacity can be maintained at reduced computation by replacing the feed-forward layer in each transformer layer with a mixture-of-experts (MoE) 2. The model trained on filtered data shows consistently better performances on both NLG and NLU tasks, where the effect of filtering is more significant on the former tasks 3. Filtered pretraining corpora play a crucial role in the generation capability of LLMs, especially for the downstream tasks 4. The scaling of GLaM MoE models can be achieved by increasing the size or number of experts in the MoE layer. Given a fixed budget of computation, more experts contribute to a better performance
@@
1. The model can be fine-tuned to learn to call different external information resources and tools